OpenStudy (anonymous):

The Celsius temperature C is given in terms of the Fahrenheit temperature F by C=5/9(F-32). Find the rate of change of C with respect to F and the rate of change of F with respect to C.

OpenStudy (anonymous):

take the derivative : dc /df = 5/9 F = (9C/5) + 32 df/dc = 9/5

OpenStudy (anonymous):

c'=5/9 f'=9/5